# Logic Distillation: Learning from Code Function by Function for Planning and Decision-making

### v1

@@ Abstract
Large language models (LLMs) have garnered increasing attention owing to their powerful logical reasoning capabilities. Generally, larger LLMs (L-LLMs) that require paid interfaces exhibit significantly superior performance compared to smaller LLMs (S-LLMs) that can be deployed on a variety of devices. Knowledge distillation (KD) aims to empower S-LLMs with the capabilities of L-LLMs, while S-LLMs merely mimic the outputs of L-LLMs, failing to get the powerful logical reasoning capabilities. Consequently, S-LLMs are helpless when it comes to planning and decision-making tasks that require logical reasoning capabilities. To tackle the identified challenges, we propose a novel framework called Logic Distillation (LD). Initially, LD employs L-LLMs to instantiate complex instructions into discrete functions and illustrates their usage to establish a function base. Subsequently, based on the function base, LD fine-tunes S-LLMs to learn the logic employed by L-LLMs in planning and decision-making. During testing, LD utilizes a retriever to identify the top-$K$ relevant functions based on instructions and current states, which will be selected and invoked by S-LLMs. Ultimately, S-LLMs yield planning and decision-making outcomes, function by function. Relevant experiments demonstrate that with the assistance of LD, S-LLMs can achieve outstanding results in planning and decision-making tasks, comparable to, or even surpassing, those of L-LLMs.
Introduction
Figure 1: The planning and decision-making outcome of one step in the pursuit game. The game consists of a pursuit scenario in which blue dots, controlled by different LLMs, aim to capture orange dots that are governed by GLM-4.
@@ Introduction
Large language models (LLMs) like GPT-3.5 (Ouyang et al. 2022) and GLM-4 (GLM et al. 2024) have been extensively applied owing to their robust capabilities in logical reasoning. Particularly, LLMs demonstrate superior performance in autonomous embodied agents, showcasing advanced planning and decision-making capabilities grounded in logical reasoning (Xi et al. 2023).
@@
Despite the remarkable capabilities of LLMs such as GPT-3.5 and GLM-4, their substantial computational requirements render them impractical for deployment on most devices (Chen et al. 2024b, a). On the other hand, numerous companies have attempted to develop relatively smaller open-source LLMs, including GLM-4-9B (GLM et al. 2024) and LLaMA-7B (Touvron et al. 2023), which are compatible with consumer-grade GPUs like the RTX 4090 Ti. In this paper, we refer to LLMs that require invocation through a paid interface as Larger-LLMs (L-LLMs), in contrast to Smaller-LLMs (S-LLMs) deployable on consumer-grade GPUs. Generally, L-LLMs exhibit significantly superior performance across various domains, particularly in logical reasoning. Nonetheless, S-LLMs have garnered extensive attention owing to their convenient deployment and cost-free nature. Consequently, an increasing number of researchers are focusing on Knowledge Distillation (KD) of LLMs (Gou et al. 2021), where L-LLMs act as teachers imparting knowledge, while S-LLMs serve as students, mimicking the outputs of teachers (Xu et al. 2024).
@@
While KD has been demonstrated to effectively enhance the capabilities of S-LLMs in numerous tasks, such as Natural Language Understanding (Dai et al. 2023; Gilardi, Alizadeh, and Kubli 2023), endowing S-LLMs with the planning and decision-making capabilities of L-LLMs through KD remains a challenge. We present the planning and decision-making outcome of a single step in a pursuit game, where LLMs control three blue dots in pursuit of an orange dot. Figure 1(a) displays the initial positions of four dots. Meanwhile, in Figure 1(b), it is evident that the L-LLM (GLM-4) effectively comprehends the game rules, enabling informed planning and decision-making based on the position of the orange dot at each step. In contrast, (c) reveals that the S-LLM (GLM-4-9B) lacks proficiency in following rules as the points enclosed by the red dashed line violate the rules of movement. Additionally, (d) highlights the issue with KD, where the student model simply mimics the output of the teacher model without comprehending the logic behind the teacher model’s planning and decision-making. Specifically, in (d), the orange dot moved one unit to the right, while the point enclosed by the purple dashed line moved one unit to the left. This stems from the KD process, where the S-LLM remembers the outputs the L-LLM has had in the past. Nevertheless, (d) demonstrates that the S-LLM is incapable of handling unseen circumstances, and fails to grasp the overall task objectives and rules. Based on the aforementioned analysis, we summarize the limitations of S-LLMs with KD as follows: 1. Ineffectiveness in following complex instructions: Despite extensive fine-tuning, S-LLMs continue to struggle with understanding intricate rules in planning and decision-making tasks. 2. Failure to comprehend the logic of L-LLMs: Fine-tuned S-LLMs simply mimic the outputs of L-LLMs and lose their planning and decision-making capabilities when encountering unknown scenarios.
@@
Recently, there has been considerable work aimed at establishing the translation between natural language and code, where a sentence is a logical code line (Feng et al. 2020; Chen et al. 2021). Drawing inspiration from this, we suggest breaking down the planning and decision-making logic of L-LLMs into multiple stages, with each stage being represented by a specific code function. Subsequently, S-LLMs engage in planning and decision-making by learning and applying the pertinent functions. More specifically, we propose Logic Distillation (LD). First, we leverage the powerful comprehension and logical reasoning capabilities of L-LLMs to decompose the rules governing planning and decision-making tasks into multiple stages, forming different functions. Subsequently, we integrate these functions along with their comments, usage examples, etc., into a function base, and use it to fine-tune S-LLMs to enhance their ability to invoke relevant functions. Besides, to alleviate the issue of S-LLMs struggling to follow complex instructions, we propose transforming generation into selection. During each stage of planning and decision-making, LD employs a retriever to provide top-$K$ relevant functions based on the current state and instructions, enabling S-LLMs to select and execute functions from this set.
@@
As depicted in Figure 2, both KD and LD require guidance from a teacher with superior capabilities. KD emphasizes the imparting of content, which involves students mimicking the teacher’s output. In contrast, the proposed LD concentrates on the underlying logic of task execution. The teacher decomposes a complex task into multiple basic logics, represented by functions, enabling the student to plan and make decision function by function.
@@
  • We analyzed the issues of Knowledge Distillation in the context of planning and decision-making in the LLMs era.
  • We propose Logic Distillation to enable S-LLMs to accomplish complex planning and decision-making tasks akin to L-LLMs.
  • We conducted experiments in different simulation scenarios to validate the effectiveness of the proposed method.
@@
Figure 3: Illustration of the proposed Logic Distillation (LD). LD consists of three components: L-LLMs, retriever, and S-LLMs. L-LLMs are responsible for decomposing human-provided rules and instantiating them as basic functions to construct a function base. Besides, L-LLMs offer a user manual that explains the usage of these functions, including details such as rule descriptions and code comments. The retriever is in charge of retrieving the top-$K$ functions based on the insturctions and states. S-LLMs select the appropriate functions for different stages of the task. Subsequently, S-LLM will systematically plan and make decisions function by function.
Related Work
@@
Large language models (LLMs) (Chowdhery et al. 2022; Thoppilan et al. 2022; Zhao et al. 2023; Koundinya Gundavarapu et al. 2024) are trained on broad data and can be easily adapted to a wide range of tasks (Bommasani et al. 2021), which have been applied to education (Biswas 2023; Kasneci et al. 2023), healthcare (Thirunavukarasu et al. 2023; Peng et al. 2023; Guo et al. 2024), finance (Wu et al. 2023), etc. However, LLMs with impressive capabilities often suffer from size limitations, making them impractical to run on various devices and costly for invoking their interfaces (Chen et al. 2024b). In this paper, we refer to such LLMs as larger LLMs (L-LLMs). In contrast, we refer to LLMs with a smaller parameter scale that can be deployed on various devices as smaller LLMs (S-LLMs). Generally, S-LLMs are more flexible than L-LLMs, but their performance is significantly inferior to that of L-LLMs.
@@
In order to endow S-LLMs with the superior capabilities of L-LLMs, Knowledge Distillation (KD) of LLMs has become a focal point of related research (Xu et al. 2024). In this context, L-LLMs like GPT-4 or GLM-4 are highly skilled teachers, while S-LLMs are students that learn to mimic the outputs of teachers (He et al. 2023; Gu et al. 2024; Agarwal et al. 2024; Liu et al. 2023). Besides, numerous works have also focused on the reasoning steps of LLMs. Distilling step-by-step (Hsieh et al. 2023) extracts L-LLM rationales as additional supervision for training S-LLMs within a multi-task framework. The Orca framework (Mukherjee et al. 2023) augments the prompt-response data pairs by incorporating a system message designed to facilitate student models’ comprehension of the reasoning process. Subsequently, Orca 2 (Mitra et al. 2023) advances this approach by training the student model to discern the most efficacious solution strategy for individual tasks, guided by the performance metrics established by Orca. However, these methods are essentially still making S-LLMs mimic the outputs of L-LLMs, rather than comprehending the reasoning logic.
Methodology
@@
This paper investigates interactive planning and decision-making tasks, where interactions can be divided into multiple steps (each interaction counts as one step), and each step can be further divided into multiple stages (several stages collectively complete one planning and decision-making process).
@@
As illustrated in Figure 3, we explore the Logic Distillation (LD) to empower S-LLMs to engage in planning and decision-making akin to L-LLMs. LD first decomposes rules (instructions) $x$ into multiple stages with L-LLMs and converts the execution logic into the corresponding code functions (such as calculating the distance between points). Then, LD will construct a function base containing functions and user manual of these functions. Besides, LD employs a retriever to find the top-$K$ functions that most relevant to the current stage and $x$. Subsequently, S-LLMs will select functions $f$ one by one to plan and make decisions. Overall, in this context, LD consists of three components: (i) L-LLMs, (ii) retriever, (iii) S-LLMs.
L-LLMs
@@
L-LLMs exhibit exceptional capabilities in planning and decision-making. To distill these capabilities into the S-LLMs, we propose instantiating the reasoning logic of L-LLMs through functions:
$p_{\theta_{L}}(f,u|x)=\prod_{i}p_{\theta_{L}}(y_{i}|x,y_{1:i-1}),f,u=\prod_{i}y_{i}$ (1)
@@
where L-LLMs $p_{\theta_{L}}$ parametrized by $\theta_{L}$ that generates a current token $y_{i}$ based on a context of the previous $i-1$ tokens, multiple $y$ constitute a function $f$ and user manual (includes rule explanations, code comments, corresponding invocation stages, and so on). In addition, a collection of $f$ and $u$ forms the function base $D_{f}$.
Retriever
@@
For the retriever $p_{\theta_{R}}(f|x)$, we have proposed two solutions tailored for function base of different scales. When the scale of the function base is large, we follow prior work (Lewis et al. 2020) to implement retrieval component based on DPR (Karpukhin et al. 2020), and retriever $p_{\theta_{R}}(f|x)$ follows:
$p_{\theta_{R}}(f|x)\propto exp(d(f)^{\top},q(x))\rightarrow f_{candidates}=[f_{1},\cdots,f_{K}]$ (2)
@@
where $d(f)$ is a dense representation of functions (including code comments and rule descriptions), and $q(x)$ is a query representation. Retriever $p_{\theta_{R}}(f|x)$ will return a top-$K$ list, where the $K$ functions with highest prior probability. As for small-scale function base, we will fine-tune S-LLMs so that S-LLMs can directly select and utilize the appropriate functions from base $D_{f}$ to plan and make decisions.
S-LLMs
@@
We fine-tune S-LLMs to enable them to comprehend the functionality and the appropriate invocation timing of different functions. S-LLMs first select a function $f_{j}$ from $D_{f}$ or $f_{candidates}$ for stage $j$:
$p_{\theta_{S}}(f_{j}|x,u)=max([p_{\theta_{S}}(f_{1}|x,u),\cdots,p_{\theta_{S}}(f_{K}|x,u)])$ (3)
@@
Then, function $f_{j}$ will be executed to obtain the intermediate result of the $j$-th stage.
$o_{j}=f_{j}(o_{j-1}),\quad o_{0}=x$ (4)
@@
If there are $J$ stages in a step of the task, then the S-LLMs will select $J$ functions, and the planning and decision-making outputs for that step will be:
$O=o_{J}=f_{J}(x,o_{J-1})$ (5)
@@
If $O$ meets the requirements of the task, the planning and decision-making process will be halted. Otherwise, the $O$ will be regarded as input for the next step. The proposed LD is summarized in Algorithm 1.
Emergency handling of S-LLMs in LD
@@
An advantage of LLMs is their ability to respond to various situations. When using LLMs to control embodied agents, they often can respond to unforeseen circumstances. For instance, when LLMs control unmanned vessels for maritime exploration, they might navigate from the open sea to archipelagos, and LLMs can analyze the specific terrain, enabling swift traversal of the archipelago.
@@
Typically, emergency situations (such as avoiding whirlpools when controlling unmanned vessels) are simpler compared to the initial various instructions and rules. Therefore, in the LD framework, S-LLMs can be used to transform emergency $x_{E}$ into functions and add them to the function candidate list $f_{candidates}$:
$p_{\theta_{S}}(f_{E},u|x)=\prod_{i}p_{\theta_{S}}(y_{i}|x_{E},y_{1:i-1}),f_{E},u=\prod_{i}y_{i}$ (6)
@@
It should be noted that, in contrast to KD, which necessitates S-LLMs to memorize massive L-LLMs’ outputs, LD merely requires S-LLMs to remember the usage rules of functions. Consequently, LD preserves more general capabilities of LLMs, including function generation.
Algorithm 1 Testing of Logic Distillation
@@
Output: the planning and decision-making outcome $[o_{1},o_{2},\cdots]$
@@
2:  Building function base $D_{f}$ by $f$ and $u$.
3:  Initialize $O$.
4:  while Planning and decision-making output $O$ of one step does not meet the task requirements do
@@
6:        Retrieve top-$K$ functions $[f_{1},\cdots,f_{K}]$ with $p_{\theta_{R}}$ and $x$ by Equation 2.
7:        S-LLMs select the most suitable function $f_{j}$ from $[f_{1},\cdots,f_{K}]$ for stage $j$.
@@
12:        Generate functions $f_{E}$ by Equation 6 and add $f_{E}$ into $[f_{1},\cdots,f_{K}]$.

### v2

# Logic Distillation: Learning from Code Function by Function for Decision-making Tasks

@@ Abstract
Large language models (LLMs) have garnered increasing attention owing to their powerful comprehension and generation capabilities. Generally, larger LLMs (L-LLMs) that require paid interfaces exhibit significantly superior performance compared to smaller LLMs (S-LLMs) that can be deployed on a variety of devices. Knowledge distillation (KD) aims to empower S-LLMs with the capabilities of L-LLMs, while S-LLMs merely mimic the outputs of L-LLMs, failing to get the powerful decision-making capability for new situations. Consequently, S-LLMs are helpless when it comes to continuous decision-making tasks that require logical reasoning. To tackle the identified challenges, we propose a novel framework called Logic Distillation (LD). Initially, LD employs L-LLMs to instantiate complex instructions into discrete functions and illustrates their usage to establish a function base. Subsequently, LD fine-tunes S-LLMs based on the function base to learn the logic employed by L-LLMs in decision-making. During testing, S-LLMs will yield decision-making outcomes, function by function, based on current states. Experiments demonstrate that with the assistance of LD, S-LLMs can achieve outstanding results in continuous decision-making tasks, comparable to, or even surpassing, those of L-LLMs. The code and data for the proposed method are provided for research purposes https://github.com/Anfeather/Logic-Distillation.
1 Introduction
Figure 1: The outcome of one step in the pursuit game.
@@ 1 Introduction
Large language models (LLMs) Ouyang et al. (2022), such as GPT-4 Achiam et al. (2023) and GLM-4 GLM et al. (2024), have been extensively applied owing to their powerful capabilities like comprehension and generation. Particularly, LLMs demonstrate superior performance in autonomous embodied agents, showcasing advanced decision-making capabilities grounded in comprehending instructions and logical reasoning Xi et al. (2023).
@@
Despite the remarkable capabilities of LLMs, such as GPT-4 and GLM-4, their substantial computational requirements render them impractical for deployment on most devices Chen et al. (2024b, a, 2025). On the other hand, numerous companies have attempted to develop relatively smaller open-source LLMs, including GLM-4-9B GLM et al. (2024) and LLaMA-7B Touvron et al. (2023), which are compatible with consumer-grade GPUs like RTX 3090 Ti. In this paper, we refer to LLMs that cannot be deployed on most devices and require invocation through a paid interface as larger LLMs (L-LLMs), in contrast to smaller LLMs (S-LLMs) deployable on consumer-grade GPUs. Generally, L-LLMs exhibit significantly superior performance across various domains, particularly in logical reasoning. Nonetheless, S-LLMs have garnered extensive attention owing to their convenient deployment and cost-free nature. Consequently, an increasing number of researchers are focusing on Knowledge Distillation (KD) of LLMs Gou et al. (2021), where L-LLMs act as teachers imparting knowledge, while S-LLMs serve as students, mimicking the outputs of teachers Xu et al. (2024).
@@
While KD has been demonstrated to effectively enhance the capabilities of S-LLMs in numerous tasks Dai et al. (2023), endowing S-LLMs with the decision-making capability of L-LLMs continues to pose a significant challenge. As shown in Figure 1, we present the decision-making outcome of a single step in a pursuit game, where LLMs control three blue dots chase an orange dot. Figure 1(a) displays the initial positions of four dots. Meanwhile, in Figure 1(b), the L-LLM (GLM-4) effectively comprehends the game rules, enabling informed decision-making based on the position of the orange dot. In contrast, Figure 1(c) reveals that the S-LLM (GLM-4-9B) lacks proficiency in following rules as the points enclosed by the red dashed line violate the rules of movement. Additionally, Figure 1(d) highlights the issue with KD, where the student merely mimics the output of the teacher without comprehending the logic behind the teacher’s decision-making. Specifically, in Figure 1(d), the orange dot moved one unit to the right, while the point enclosed by the purple dashed line moved one unit to the left. Such results stem from the KD process, where the S-LLM remembers the outputs the L-LLM has had in the past. Based on the aforementioned analysis, we summarize the limitations of S-LLMs and KD in decision-making tasks as follows: 1. Ineffectiveness in following complex instructions: Despite extensive fine-tuning, S-LLMs continue to struggle with following intricate rules in decision-making tasks. 2. Failure to comprehend the logic of L-LLMs: Fine-tuned S-LLMs merely mimic the outputs of L-LLMs and lose their decision-making capabilities when encountering unknown scenarios.
@@
Recently, there has been considerable work aimed at establishing the translation between natural language and code, where a sentence is a logical code line Feng et al. (2020); Chen et al. (2021). Drawing inspiration from this, we suggest breaking down the decision-making logic of L-LLMs into multiple stages, with each stage being represented by a specific code function. Subsequently, S-LLMs engage in decision-making by learning and applying the pertinent functions. More specifically, we propose Logic Distillation (LD). First, we leverage the powerful comprehension and logical reasoning capabilities of L-LLMs to decompose the rules governing decision-making tasks into multiple stages, forming different functions. Subsequently, we integrate these functions along with their comments, usage examples, etc., into a function base, and use it to fine-tune S-LLMs to enhance their ability to invoke relevant functions. Besides, to alleviate the issue of S-LLMs struggling to follow complex instructions, we propose transforming generation into selection. During each stage of decision-making, S-LLMs will select and invoke functions based on the current state.
@@
As depicted in Figure 2, both KD and LD require guidance from a teacher with superior capabilities. KD emphasizes the imparting of content, which involves students mimicking the teacher’s output. In contrast, the proposed LD concentrates on the underlying logic of task execution. The teacher decomposes a complex task into multiple basic logics, represented by functions, enabling the student to make decision function by function.
@@
  • We analyzed the issues of Knowledge Distillation in the context of decision-making in the LLMs era.
  • We propose Logic Distillation to enable S-LLMs to accomplish decision-making tasks akin to L-LLMs.
  • We conducted experiments in different scenarios to validate the effectiveness of the proposed method.
@@
Figure 3: Illustration of the proposed Logic Distillation (LD). LD consists of three components: L-LLMs, retriever, and S-LLMs. L-LLMs are responsible for decomposing human-provided rules and instantiating them as basic functions to construct a function base. Besides, L-LLMs offer a user manual that explains the usage of these functions, including details such as rule descriptions and code comments. The retriever is in charge of retrieving the top-$K$ functions based on the insturctions and states. S-LLMs select the appropriate functions for different stages of the task. Subsequently, S-LLMs will systematically make decisions function by function.
2 Related Work
@@
Large language models (LLMs) Ouyang et al. (2022); Chowdhery et al. (2022); Thoppilan et al. (2022); Zhao et al. (2023); Koundinya Gundavarapu et al. (2024) are trained on broad data and can be easily adapted to a wide range of tasks Bommasani et al. (2021), which have been applied to education Biswas (2023); Kasneci et al. (2023), healthcare Thirunavukarasu et al. (2023); Peng et al. (2023); Guo et al. (2024), finance Wu et al. (2023), etc. However, LLMs with impressive capabilities often suffer from size limitations, making them impractical to run on most devices and costly for invoking their interfaces Chen et al. (2024b).
@@
In order to endow S-LLMs with the superior capabilities of L-LLMs, Knowledge Distillation (KD) of LLMs has become a focal point of related research Xu et al. (2024). In this context, L-LLMs like GPT-4 or GLM-4 are highly skilled teachers, while S-LLMs are students that learn to mimic the outputs of teachers He et al. (2023); Gu et al. (2024); Agarwal et al. (2024); Liu et al. (2023). Besides, numerous works have also focused on the reasoning steps of LLMs. Distilling step-by-step Hsieh et al. (2023) extracts L-LLM rationales as additional supervision for training S-LLMs within a multi-task framework. The Orca framework Mukherjee et al. (2023) augments the prompt-response data pairs by incorporating a system message designed to facilitate student models’ comprehension of the reasoning process. Subsequently, Orca 2 Mitra et al. (2023) advances this approach by training the student model to discern the most efficacious solution strategy for individual tasks, guided by the performance metrics established by Orca. However, these methods are essentially still making S-LLMs mimic the outputs of L-LLMs, rather than comprehending the reasoning logic.
3 Methodology
@@
This paper investigates interactive decision-making tasks, where interactions can be divided into multiple steps (each interaction counts as one step), and each step can be further divided into multiple stages (several stages collectively complete one decision-making process).
@@
As illustrated in Figure 3, we explore the Logic Distillation (LD) to empower S-LLMs to engage in decision-making akin to L-LLMs. LD first decomposes rules (instructions) $x$ into multiple stages with L-LLMs and converts the decision-making logic into the corresponding code functions (such as calculating the distance between points). Then, LD will construct a function base containing functions and user manual of these functions. Besides, LD employs a retriever to find the top-$K$ functions that most relevant to the current states and $x$. Subsequently, S-LLMs will select functions $f$ one by one to make decisions. Overall, LD consists of three components: (i) L-LLMs, (ii) retriever, (iii) S-LLMs.
3.1 L-LLMs
@@
L-LLMs exhibit exceptional capabilities in decision-making. To distill these capabilities into the S-LLMs, we propose instantiating the logic of L-LLMs through functions:
$p_{\theta_{L}}(f,u|x)=\prod_{i}p_{\theta_{L}}(y_{i}|x,y_{1:i-1})$ (1)
@@
where L-LLMs $p_{\theta_{L}}$ parametrized by $\theta_{L}$ that generates a current token $y_{i}$ based on a context of the previous i1i-1 tokens, multiple $y$ constitute a function $f$ and user manual (includes rule explanations, code comments, corresponding invocation stages, and so on). In addition, a collection of $f$ and $u$ forms the function base $D_{f}$.
3.2 Retriever
@@
For the retriever $p_{\theta_{R}}(f|x,s)$, we have proposed two solutions tailored for function base of different scales. When the scale of the function base is large, we follow prior work Lewis et al. (2020) to implement retrieval component based on DPR Karpukhin et al. (2020), and retriever $p_{\theta_{R}}(f|x,s)$ follows:
$p_{\theta_{R}}(f|x,s)\propto exp(d(f)^{\top},q(x,s))\rightarrow f_{c}=[f_{1},\cdots,f_{K}]$ (2)
@@
where $s$ is current states, $d(f)$ is a dense representation of functions (including code comments and rule descriptions), and $q(x,s)$ is a query representation. Retriever $p_{\theta_{R}}(f|x,s)$ will return a top-$K$ list $f_{c}$, where the $K$ functions with highest prior probability. As for small-scale function base, we will fine-tune S-LLMs so that S-LLMs can directly select and utilize the appropriate functions from base $D_{f}$ to make decisions.
3.3 S-LLMs
@@
We fine-tune S-LLMs to enable them to comprehend the functionality and the appropriate invocation timing of different functions. S-LLMs first select a function $f_{j}$ from $D_{f}$ or $f_{c}$ for stage $j$:
$p_{\theta_{S}}(f_{j}|x,s)=max([p_{\theta_{S}}(f_{1}|x,s),\cdots,p_{\theta_{S}}(f_{K}|x,s)])$ (3)
@@
Then, function $f_{j}$ will be executed to obtain the intermediate result of the $j$-th stage:
$o_{j}=f_{j}(x,s),\quad s=o_{j-1}$ (4)
@@
If there are $J$ stages in a step of the task, S-LLMs will select $J$ functions, and the decision-making outputs for that step will be:
$O=o_{J}=f_{J}(x,s),\quad s=o_{J-1}$ (5)
@@
If $O$ meets the requirements of the task, the decision-making process will be halted. Otherwise, $O$ will be regarded as input for the next step.
3.4 Emergency handling of S-LLMs in LD
@@
An advantage of LLMs is their ability to respond to various situations. When using LLMs to control embodied agents, they can respond to unforeseen circumstances. For instance, when LLMs control unmanned vessels for maritime exploration, they might navigate from the open sea to archipelagos, and LLMs can analyze the specific terrain, enabling swift traversal of the archipelago.
@@
Typically, actions required in an emergency situation (such as avoiding whirlpools when controlling unmanned vessels) are simpler compared to the initial various instructions and rules. Therefore, in the LD framework, S-LLMs will transform emergency $x_{E}$ into functions and add them to the function candidate list $f_{c}$:
$p_{\theta_{S}}(f_{E},u|x_{E},s)=\prod_{i}p_{\theta_{S}}(y_{i}|x_{E},s,y_{1:i-1}),f_{E},u=\prod_{i}y_{i}$ (6)
@@
It should be noted that, in contrast to KD, which necessitates S-LLMs to memorize massive L-LLMs’ outputs, LD merely requires S-LLMs to remember the usage of functions. Consequently, LD preserves more general capabilities of LLMs, including function generation. The proposed LD is summarized in Algorithm 1.
Algorithm 1 Logic Distillation
@@
Output: the decision-making outcome $o_{1},o_{2},\cdots]$ .
@@
2: Building function base $D_{f}$ with $f$ and $u$.
3: Initialize $O$, $s$.
4:while Decision-making output $O$ of one step does not meet the task requirements do
@@
6:   Retrieve top-$K$ functions $f_{1},\cdots,f_{K}]$ with $p_{\theta_{R}}$, $x$ and $s$ by Equation 2.
7:   S-LLMs select the most suitable function $f_{j}$ from $f_{1},\cdots,f_{K}]$ for stage $j$.
@@
12:   Generate functions $f_{E}$ by Equation 6 and add $f_{E}$ into $f_{1},\cdots,f_{K}]$.